Find the modulus and the principal argument of each of the given complex numbers.
(a) 3 − 4i,

(b) −2 + i,

(c) 1/1 + i √3

(d) (7 − i)/(-4 - 3i)

(e) 5(cos π/3 + isin π/3)

(f) cos 2π/3 − sin 2π/3.
1
Expert's answer
2020-06-03T19:01:23-0400

The formula for modulus (r) and principal argument (θ) of a complex number (z = x + iy) is :


r = ∣z∣ = "\sqrt{\smash[b]{x² + y²}}"

θ = arctan(y/x)


(a) z = 3 - 4i

Modulus, r = "\sqrt{\smash[b]{3² + (-4)²}}"

​ r = "\sqrt{\smash[b]{9 + 16}}" = "\sqrt{\smash[b]{25}}"

∴ r = 5


Principal argument, θ = arctan(-4/3)

θ = -53.13°

(b) z = -2 + i

Modulus, r = "\sqrt{\smash[b]{(-2)² + (1)²}}"

r = "\sqrt{\smash[b]{4 + 1}}"

∴ r = "\sqrt{\smash[b]{5}}"


Principal argument, θ = arctan(1/(-2))

θ=−26.56°

Since the given complex number lies in the second quadrant,

Therefore, θ=(−26.56+180)°

θ=153.44°


(c) z = "1/(1 + i\sqrt{\smash[b]{3}})"


"z = (1 - i\sqrt{\smash[b]{3}})/((1 + i\sqrt{\smash[b]{3}})(1 - i\sqrt{\smash[b]{3}}))"


"z = (1 - i\sqrt{\smash[b]{3}})/(1² - (i\sqrt{\smash[b]{3}})²)"


"z = (1 - i\sqrt{\smash[b]{3}})/(1 + 9)"


"z = (1 - i\sqrt{\smash[b]{3}})/10"


"z = (1/10) - i(\sqrt{\smash[b]{3}}/10)"



Modulus, r ="\sqrt{\smash[b]{(1/10)² + (\sqrt{\smash[b]{3}}/10)²}}"

"r = \sqrt{\smash[b]{(4/100)}}"


∴ r = (2/10)


Principal argument, θ = arctan(-"(\sqrt{\smash[b]{3}}/10)/(1/10)")


θ=arctan"(- \sqrt{\smash[b]{3}})"

θ=−(π/3)

θ=−60°


(d) z = (7 − i)/(-4 - 3i)

z = ((7 - i)(-4 + 3i))/((-4 - 3i)(-4 + 3i))

z = (-28 + 21i +4i + 3)/((-4)² - (3i)²)

z = (-25 + 25i)/(16 + 9)

z = (-25 + 25i)/25

z = -1 + i


Modulus, r ="\sqrt{\smash[b]{(-1)² + 1²}}"

r ="\sqrt{\smash[b]{1 + 1}}"


r="\sqrt{\smash[b]{2}}"


Principal argument, θ = arctan(1/(-1))

θ = arctan(-1)

θ = -45°

Since the given complex number lies in the second quadrant,

Therefore, θ=(−45+180)°

θ=135°


(e) z = 5(cos(π/3)+isin(π/3))

z= 5cos(π/3)+i5sin(π/3)


Modulus, r ="\sqrt{\smash[b]{(5cos(π/3))² + (5sin(π/3))²}}"

r = "\sqrt{\smash[b]{5²(cos²(π/3)+sin²(π/3))}}"

r = 5 "\sqrt{\smash[b]{1}}"


∴ r = 5


Principal argument, θ = arctan(5sin(π/3)/5cos(π/3))

θ = arctan(tan(π/3))

θ = π/3

θ = 60°


(f) z = cos 2π/3 − sin 2π/3

z = (-1/2) -"(\sqrt{\smash[b]{3}}/2)"


"z = - (1 + \sqrt{\smash[b]{3}})/2"


Modulus, r ="\sqrt{\smash[b]{(cos(2π/3) - sin(2π/3))² + 0}}"


"r = \sqrt{\smash[b]{(cos²(2π/3) + sin²(2π/3) - 2sin(2π/3)cos(2π/3)}}"


"r = \sqrt{\smash[b]{(1 - sin(4π/3))}}"


"r = \sqrt{\smash[b]{(1 + (\sqrt{\smash[b]{3}}/2)) }}"


∴ r ="\sqrt{\smash[b]{(2 + \sqrt{\smash[b]{3}})/2) }}"


Principal argument, θ = arctan(0/(cos(2π/3) - sin(2π/3)))

θ=arctan(0)

Since the given complex number lies in the second quadrant,

Therefore, θ=(0+180)°

θ=180°